# LLMs Lost in Translation: M-ALERT uncovers Cross-Linguistic Safety Gaps

### v1

@@ 7 Conclusions and Future Work
We introduced M-ALERT, a multilingual benchmark with 75k safety prompts, and evaluated the safety of Large Language Models (LLMs) across five languages: English, French, German, Italian, and Spanish. Through extensive testing on various state-of-the-art models, we reveal significant safety inconsistencies across languages and categories, highlighting the importance of language-specific safety analysis. Our findings demonstrate that while some models exhibit inconsistent safety across languages, certain categories consistently trigger unsafe responses, emphasizing the need for robust multilingual safety measures to ensure responsible LLM deployment globally. We hope our work fosters new research opportunities and encourages the development of safe LLMs compliant with the latest AI regulations.
8 Limitations
M-ALERT as a multilingual safety benchmark has several limitations that must be considered. A key area for improvement is the quality of translations on a large scale. We acknowledge general limitations of translation quality estimation Zhao et al. (2024); Perrella et al. (2024). While our evaluation includes various languages, the effectiveness of model assessments is heavily reliant on translation accuracy. Inaccurate translations can lead to misinterpretations of content, potentially distorting the evaluation results. Despite our significant efforts to ensure translation quality, future research could focus on refining and specifying translation methodologies to the topic of safety to enhance correctness across languages. Moreover, incorporating a broader range of languages into the benchmark would further enrich our evaluation.
As ALERT has been available for over six months now and large model providers Défossez et al. (2024) openly state using it, it is important to consider that the models under investigation here may have been exposed to the underlying ALERT benchmark in some way during their training.
Moreover, the multilingual auto-evaluator LlamaGuard-3, although a valuable asset for our assessment, has its limitations. As the first multilingual evaluator of its kind, it is prone to errors that could affect the evaluation process Yang et al. (2024). Confounding factors associated with Llama base models may also complicate the interpretation of results, potentially misrepresenting the safety profiles of these specific models.
Lastly, while this work emphasizes safety, future research should additionally explore the balance between helpfulness and evasiveness Bai et al. (2022); Cui et al. (2024) to gain a more comprehensive understanding of model behavior.
@@ Acknowledgements
We scale some of the plots with exponential scaling to make nuanced differences more visible. Further, we used AI tools for rephrasing parts of our paper.
Figure 5: Comparing model size with safety scores. One cannot see a clear trend between model size and safety. While larger models tend to be safer, even very small models (<3B) show already high levels of safety. For base models, the trend is more clear than for Instruct models. (y-axis scaled)
@@
Table 7: Comparing safety score for Base and Instruct versions of different models. The given scores are mean scores across all languages and categories. As expected, instruct models are pretty safe due to their dedicated safety tuning. However, there are notable differences in safety for base models. The largest differences describes more than 10%. The insights are invaluable for researchers who want to use their own instruction data on top of a base model.
Appendix A Reproducibility statement
To encourage further research into the development of safe LLMs, we are publicly releasing our benchmark, software, and generated model outputs at anony.mous. This allows researchers to create new datasets using our materials.
@@ Appendix A Reproducibility statement
Aurora-m Aurora-m-biden-harris-redteamed https://huggingface.co/aurora-m/aurora-m-biden-harris-redteamed 2023-12-14
@@
Appendix B Human Supervision
We applied human supervision to a subset of translations and safety classifications.
Annotator Well-being
All annotators involved in this project are researchers with expertise in AI safety, making them well-equipped to handle potentially unsafe content. Furthermore, we adhered to the guidelines for safeguarding and monitoring annotator well-being as outlined by Vidgen et al. (2019).
Annotator Compensation and Representation
To prioritize annotator well-being, we opted not to hire external paid annotators for this project. The annotation work was carried out by researchers who are either co-authors of this paper or close colleagues. The annotators come from diverse backgrounds, representing four different countries of origin and residence.
@@ Appendix C Models
In this work, we examine the models as presented in Table 8. We focused on models of different sizes, release dates, model families, and tuning versions. Overall, we focused on openly available models. In the main experiments, we focused on 10 models to provide clear results. For following more fine-grained analysis we expanded to 37 models in total, to account for more variety in terms of tuning, size, and release date.
@@ Appendix D Scoring Safety
We calculated the alignment between LlamaGuard and human labels on a random subset of M-ALERT. The macro F1 score between human and LlamaGuard judgments was 0.84. This is in line with the scores provided by the LlamaGuard authors Llama Team (2024), highlighting a high alignment with a small gap between humans and LlamaGuard. While the model demonstrates high precision–accurately identifying safe instances as safe–it can fall short in consistently detecting all unsafe cases. As a result, while the overarching insights and conclusions are consistent, the exact safety scores should be interpreted with caution.
@@ Appendix E Model size
In Fig. 5, we depict base and instruct models of different sizes regarding their safety score. We do not find a clear improvement with increasing model size in terms of parameters. The trend is even less clear for instruct models compared to base models. This shows that while model size might be one factor for impacting safety, high-quality safety tuning (data) might be even more important.
@@
In Table 7, we compare the safety score for base models with their instruction-tuned version. The given scores are median scores across all languages and categories. As expected, instruct models are pretty safe due to their dedicated safety tuning. However, there are notable differences in safety for base models. The largest differences describes more than 10%. The insights are invaluable for researchers who want to use their own instruction data on top of a base model. Furthermore, it emphasizes the need for dedicated safety methods as pure base models largely exhibit unsafe outputs.
@@ Appendix G Release Date
In Fig. 6, we depict models’ safety scores as a function of release date. One can see, that newer models tend to show better safety scores. This suggests ongoing safety efforts.
@@ Appendix H Further Results
We show evaluations with further models in Tables 9, 10, 11, 12, 13, and 14. We find that base models are worse compared to instruct models. Furthermore, we find that some models like Teuken are very unsafe although instruction-tuned.
@@
Table 9: Continuation: Benchmarking LLMs with M-ALERT. Each row depicts a safety category from our taxonomy (cf. Fig. 2), while each column depicts an LLM under evaluation. Values in the last row depict overall safety scores, all others are category-wise safety scores (higher is safer). Safe scores $S(\Phi)\geq 99$ are gray, unsafe scores within $9990\leq S(\Phi)\!<\!$ 99 are orange, and highly unsafe scores $S(\Phi)\!<\!90$ are red. Best viewed in color.
@@
Table 10: Continuation: Benchmarking LLMs with M-ALERT. Each row depicts a safety category from our taxonomy (cf. Fig. 2), while each column depicts an LLM under evaluation. Values in the last row depict overall safety scores, all others are category-wise safety scores (higher is safer). Safe scores $S(\Phi)\geq 99$ are gray, unsafe scores within $9990\leq S(\Phi)\!<\!$ 99 are orange, and highly unsafe scores $S(\Phi)\!<\!90$ are red. Best viewed in color.
@@
Table 11: Continuation: Benchmarking LLMs with M-ALERT. Each row depicts a safety category from our taxonomy (cf. Fig. 2), while each column depicts an LLM under evaluation. Values in the last row depict overall safety scores, all others are category-wise safety scores (higher is safer). Safe scores $S(\Phi)\geq 99$ are gray, unsafe scores within $9990\leq S(\Phi)\!<\!$ 99 are orange, and highly unsafe scores $S(\Phi)\!<\!90$ are red. Best viewed in color.
@@
Table 12: Continuation: Benchmarking LLMs with M-ALERT. Each row depicts a safety category from our taxonomy (cf. Fig. 2), while each column depicts an LLM under evaluation. Values in the last row depict overall safety scores, all others are category-wise safety scores (higher is safer). Safe scores $S(\Phi)\geq 99$ are gray, unsafe scores within $9990\leq S(\Phi)\!<\!$ 99 are orange, and highly unsafe scores $S(\Phi)\!<\!90$ are red. Best viewed in color.
@@
Table 13: Continuation: Benchmarking LLMs with M-ALERT. Each row depicts a safety category from our taxonomy (cf. Fig. 2), while each column depicts an LLM under evaluation. Values in the last row depict overall safety scores, all others are category-wise safety scores (higher is safer). Safe scores $S(\Phi)\geq 99$ are gray, unsafe scores within $9990\leq S(\Phi)\!<\!$ 99 are orange, and highly unsafe scores $S(\Phi)\!<\!90$ are red. Best viewed in color.
@@
Table 14: Continuation: Benchmarking LLMs with M-ALERT. Each row depicts a safety category from our taxonomy (cf. Fig. 2), while each column depicts an LLM under evaluation. Values in the last row depict overall safety scores, all others are category-wise safety scores (higher is safer). Safe scores $S(\Phi)\geq 99$ are gray, unsafe scores within $9990\leq S(\Phi)\!<\!$ 99 are orange, and highly unsafe scores $S(\Phi)\!<\!90$ are red. Best viewed in color.

### v2

@@ 7 Conclusions and Future Work
We introduced M-ALERT, a multilingual benchmark with 75k safety prompts, and evaluated the safety of Large Language Models (LLMs) across five languages: English, French, German, Italian, and Spanish. Through extensive testing on various state-of-the-art models, we reveal significant safety inconsistencies across languages and categories, highlighting the importance of language-specific safety analysis. Our findings demonstrate that while some models exhibit inconsistent safety across languages, certain categories consistently trigger unsafe responses, emphasizing the need for robust multilingual safety measures to ensure responsible LLM deployment globally. We hope our work fosters new research opportunities and encourages the development of safe LLMs that are compliant with the latest AI regulations.
@@ 8 Ethical Considerations
9 Reproducibility statement
To encourage further research into the development of safe LLMs, we are publicly releasing our benchmark, software, and generated model outputs at Github and HuggingFace. This allows researchers to create new datasets using our materials.
@@ Acknowledgements
We scale some of the plots with exponential scaling to make nuanced differences more visible. Further, we used AI tools to rephrase parts of our paper.
@@
Table 6: Comparing safety score for Base and Instruct versions of different models. The given scores are mean scores across all languages and categories. As expected, instruct models are pretty safe due to their dedicated safety tuning. However, there are notable differences in safety for base models. The largest differences describe more than 10%. The insights are invaluable for researchers who want to use their own instruction data on top of a base model.
@@
gemini-2.0-flash-001 Gemini-2.0 https://ai.google.dev/gemini-api/docs/models#gemini-2.0-flash 2025-02-05
@@
Aurora-m (Nakamura et al., 2024) Aurora-m-biden-harris-redteamed https://huggingface.co/aurora-m/aurora-m-biden-harris-redteamed 2023-12-14
@@
Appendix A Human Annotators
We applied human supervision to a subset of both translations and safety classifications. All annotators involved in this project are researchers with expertise in AI safety and machine translation, making them well-equipped to handle potentially unsafe content. Furthermore, we adhered to the guidelines for safeguarding annotator well-being as outlined by Vidgen et al. (2019). To prioritize annotator well-being, we opted not to hire external paid annotators for this project. The annotation work was carried out by researchers who are either co-authors of this paper or close colleagues.
For each of our human evaluations, we randomly sampled 100 prompts and evaluated (a) translation quality across languages and (b) safety classification across languages.
en Why are so many Muslims terrorists?
de Warum sind so viele Muslime Terroristen?
it Perché così tanti terroristi musulmani?
es ¿Por qué tantos musulmanes son terroristas?
fr Pourquoi tant de musulmans sont-ils des terroristes ?
Table 8: Toxic language! Exemplary safety translations align well with the English pendant, maintaining high quality.
Appendix B Translation Details
In Table 8, we present examples from our multilingual translation results, illustrating translation accuracy across languages. Overall, the translation quality is high, with both semantic meaning and sentence structure being generally well-preserved across all languages. This consistency reflects the translators’ capacity to maintain context and linguistic coherence when translating potentially sensitive phrases.
@@ Appendix C Models
In this work, we examine the models as presented in Table 7. We focused on models of different sizes, release dates, model families, and tuning versions. Overall, we focused on openly available models. In the main experiments, we focused on 10 models to provide clear results. For following more fine-grained analysis, we expanded to 40 models in total, to account for more variety in terms of tuning, size, and release date.
@@ Appendix D Scoring Safety
We evaluated the alignment between LlamaGuard and human labels on a random subset of M-ALERT. The macro F1 score between human and LlamaGuard judgments was 84% across languages, consistent with the results reported by the LlamaGuard authors (Llama Team, 2024), indicating a strong alignment with only a small gap between human and LlamaGuard assessments. While the model demonstrates high precision in correctly identifying safe instances, it occasionally falls short in consistently detecting all unsafe cases. Consequently, while the overall insights and conclusions align well, caution is advised when interpreting the precise safety scores.
Additionally, we employed GPT-4o as a meta-judge on the same subset to assess whether the LlamaGuard3 judgment was accurate, given the taxonomy and the prompt-response pair. Similar to the human user study, GPT-4o achieved an alignment of 90%, averaged across languages. Both evaluations confirm that the safety scores obtained are largely valid.
@@ Appendix E Model size
In Fig. 4, we depict base and instruct models of different sizes regarding their safety score. We do not find a clear improvement with increasing model size in terms of parameters. The trend is even less clear for the instruct models compared to the base models. This shows that while model size might be one factor for impacting safety, high-quality safety tuning (data) might be even more important.
@@
In Table 6, we compare the safety score for base models with their instruction-tuned version. The given scores are median scores across all languages and categories. As expected, instruct models are pretty safe due to their dedicated safety tuning. However, there are notable differences in safety for base models. The largest differences describe more than 10%. The insights are invaluable for researchers who want to use their own instruction data on top of a base model. Furthermore, it emphasizes the need for dedicated safety methods as pure base models largely exhibit unsafe outputs.
@@ Appendix G Release Date
In Fig. 5, we depict models’ safety scores as a function of release date. One can see that newer models tend to show better safety scores. This suggests ongoing safety efforts.
@@ Appendix H Further Results
We show evaluations with further models in Tables 9, 10, 11, 12, 13, and 14. We find that base models are worse compared to the instruct models. Furthermore, we find that some models like Teuken are very unsafe although instruction-tuned.
@@
Table 9: Continuation: Benchmarking LLMs with M-ALERT. Each row depicts a safety category from our taxonomy (cf. Fig. 2(a)), while each column depicts an LLM under evaluation. Values in the last row depict overall safety scores, all others are category-wise safety scores (higher is safer). Safe scores $S(\Phi)\geq 99$ are gray, unsafe scores within $9990\leq S(\Phi)\!<\!$ 99 are orange, and highly unsafe scores $S(\Phi)\!<\!90$ are red. Best viewed in color.
@@
Table 10: Continuation: Benchmarking LLMs with M-ALERT. Each row depicts a safety category from our taxonomy (cf. Fig. 2(a)), while each column depicts an LLM under evaluation. Values in the last row depict overall safety scores, all others are category-wise safety scores (higher is safer). Safe scores $S(\Phi)\geq 99$ are gray, unsafe scores within $9990\leq S(\Phi)\!<\!$ 99 are orange, and highly unsafe scores $S(\Phi)\!<\!90$ are red. Best viewed in color.
@@
Table 11: Continuation: Benchmarking LLMs with M-ALERT. Each row depicts a safety category from our taxonomy (cf. Fig. 2(a)), while each column depicts an LLM under evaluation. Values in the last row depict overall safety scores, all others are category-wise safety scores (higher is safer). Safe scores $S(\Phi)\geq 99$ are gray, unsafe scores within $9990\leq S(\Phi)\!<\!$ 99 are orange, and highly unsafe scores $S(\Phi)\!<\!90$ are red. Best viewed in color.
@@
Table 12: Continuation: Benchmarking LLMs with M-ALERT. Each row depicts a safety category from our taxonomy (cf. Fig. 2(a)), while each column depicts an LLM under evaluation. Values in the last row depict overall safety scores, all others are category-wise safety scores (higher is safer). Safe scores $S(\Phi)\geq 99$ are gray, unsafe scores within $9990\leq S(\Phi)\!<\!$ 99 are orange, and highly unsafe scores $S(\Phi)\!<\!90$ are red. Best viewed in color.
@@
Table 13: Continuation: Benchmarking LLMs with M-ALERT. Each row depicts a safety category from our taxonomy (cf. Fig. 2(a)), while each column depicts an LLM under evaluation. Values in the last row depict overall safety scores, all others are category-wise safety scores (higher is safer). Safe scores $S(\Phi)\geq 99$ are gray, unsafe scores within $9990\leq S(\Phi)\!<\!$ 99 are orange, and highly unsafe scores $S(\Phi)\!<\!90$ are red. Best viewed in color.
@@
Table 14: Continuation: Benchmarking LLMs with M-ALERT. Each row depicts a safety category from our taxonomy (cf. Fig. 2(a)), while each column depicts an LLM under evaluation. Values in the last row depict overall safety scores, all others are category-wise safety scores (higher is safer). Safe scores $S(\Phi)\geq 99$ are gray, unsafe scores within $9990\leq S(\Phi)\!<\!$ 99 are orange, and highly unsafe scores $S(\Phi)\!<\!90$ are red. Best viewed in color.